OpenStudy (anonymous):

e^-x^2 . -x^2 either (-x)^2 or -(x)^2 . please look at the negative one . in bracket or out of the bracket ?? thank you.

OpenStudy (campbell_st):

\[e^{-x^2} \times -x^2 = -x^2e^{-x^2}\]

OpenStudy (campbell_st):

there is a significant difference as (-x)^2 will always be a positive value while -x^2 will always be a negative value.... not sure what the question is perhaps \[e^{-x^2 \times -x^2} = e^{x^4}\]

OpenStudy (anonymous):

hihi . sory2. act the question e^-x^2 thats all

OpenStudy (campbell_st):

\[e^{-x^2} = \frac{1}{e^{x^2}}\] perhaps this helps

OpenStudy (anonymous):

really ?

OpenStudy (campbell_st):

well using power laws negative powers are used for fractions

OpenStudy (anonymous):

owhh oke . but this is assignment question . haha. btw, thank you very much .